OpenStudy (daniellelovee):

Suppose a researcher is studying a population in which 57% of the population has a certain trait. Round your answer to the nearest tenth of a percent if needed.

OpenStudy (daniellelovee):

The researcher wants a sampling distribution with a standard deviation that is less than 6%. Would a sample size of 50 give a standard deviation that is less than 6%? How about a sample size of 100?

OpenStudy (daniellelovee):

@Directrix

Directrix (directrix):

I am still working on the other problem. Oh, well.

OpenStudy (daniellelovee):

lol yeah im sure the other problem is 9 or at least 97percent sure

Directrix (directrix):

I can't be of any help at the present because I need to review the central limit theorem. I can't find a formula that makes sense for these problems.

OpenStudy (daniellelovee):

look at what the one I gave you on the last problem

OpenStudy (daniellelovee):

alright :)

OpenStudy (daniellelovee):

@ganeshie8

ganeshie8 (ganeshie8):

\[s.d = \dfrac{\sigma}{\sqrt{n}}\]

OpenStudy (daniellelovee):

would it be 6/50?

ganeshie8 (ganeshie8):

as you can see, the standard deviation of the sampling distribution decreases quadratically as the sample size increases

OpenStudy (daniellelovee):

yes but the thing is that im not too sure about the formula that Im supposed to follow and the different percentages confuse me

OpenStudy (daniellelovee):

0.36=0.84

OpenStudy (daniellelovee):

I think im doing something wrong

ganeshie8 (ganeshie8):

You want \(s.d\) of the sampling distribution to be less than 6% of the population standard deviation. That is, \(s.d\lt 0.06\sigma\). So we can solve : \[0.06\sigma=\dfrac{\sigma}{\sqrt{n}}\]

ganeshie8 (ganeshie8):

\[\sqrt{n} = \dfrac{1}{0.06} = \dfrac{100}{6}\] square both sides and round

OpenStudy (daniellelovee):

ok so then that will be a yes for 100 sample

ganeshie8 (ganeshie8):

nope, what do you get for \(n\) when you solve

OpenStudy (daniellelovee):

50/3

OpenStudy (daniellelovee):

and you said to square root therefore 4.08

ganeshie8 (ganeshie8):

you need to square it

ganeshie8 (ganeshie8):

\(\sqrt{n} = \dfrac{50}{3}\) \(n = (\dfrac{50}{3})^2 =277.77 \)

ganeshie8 (ganeshie8):

that means, you need a sample size of at least 278 to make the standard deviation go below 6%

OpenStudy (daniellelovee):

oh then neither 50 nor 100 percent sample would work right?

ganeshie8 (ganeshie8):

they are less than 278, so none of them work

OpenStudy (daniellelovee):

ok thank you sooo much :)

ganeshie8 (ganeshie8):

yw